OpenStudy (anonymous):

is 9/13 and 18/26 porpotional yes or no

OpenStudy (lilg132):

yes

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

\[\frac{18}{26}=\frac{9*2}{13*2}= \frac{9}{13}\]